Part 1: Below is the company Information that you will use in completing Part 1. The requirements are listed after this compa7. Prepare a statement of cost of goods manufactured for January below. Use Manufacturing Overhead applied in your statement.
Answer #1
1 Predetermined overhead rate=Predicted annual overhead/Predicted annual direct material cost=30000/150000=$ 0.20 per direct material cost
2 Job 21 Job 25 Job 26 Job 27
Beginning balance 1000 3000 0 0
Current costs:
Direct materials 4000 1000 2000 3000
Direct labor 700 2000 900 1600
Applied overhead 800 200 400 600
(PDOH rate*Direct material) (0.20*4000) (0.20*1000) (0.20*2000) (0.20*3000)
Total 6500 6200 3300 5200
3 Work in process inventory
$
Beginning balance (1000+3000) 4000
Add:
Direct materials (4000+1000+2000+3000) 10000
Direct labor (700+2000+900+1600) 5200
Applied overhead (800+200+400+600) 2000
21200
Less:Cost of jobs completed
(Job 25 and Job 27) (6200+5200) 11400
Ending balance 9800
Jobs still in work in process=Job 21 and Job 26
Total cost of ending work in process for January=Total cost of Job 21 and 26=6500+3300=$ 9800
4 Finished goods inventory
$
Beginning balance 0
Add: Cost of jobs completed 11400
11400
Less: Cost of Job sold (Job 25) 6200
Ending balance 5200
Job in the finished goods at the end of January=Job completed but not sold=Job 27
Total cost of finished goods for January=Total cost of Job 27=$ 5200